OpenStudy (yoyo2443):

Use factoring and the zero-product property to solve:

OpenStudy (yoyo2443):

\[2x^2+6x=0\]

zepdrix (zepdrix):

\[\large\rm 2x\color{orangered}{x}+6\color{orangered}{x}=0\]What does each term have in common? Hmm well they certainly have x in common, I guess we can factor that out of each term,\[\large\rm \color{orangered}{x}(2x+6)=0\]What else do they have in common? Both 2 and 6 are even numbers, so they should be divisible by 2, right?

OpenStudy (yoyo2443):

Yep, and sorry for taking long to reply~

zepdrix (zepdrix):

So ya, pull a 2 out of each term,\[\large\rm 2x(x+3)=0\]the 6 loses a factor of 2, leaving a 3. The 2x loses a 2, leaving an x.

zepdrix (zepdrix):

Then apply your Zero-Factor Property: \(\large\rm 2x=0\) \(\large\rm (x+3)=0\) Setting each individual factor equal to zero. And solve for x in each equation.

OpenStudy (yoyo2443):

Ah I see, I think I get it now.